OpenStudy (anonymous):

After a discount of 10%, the price of a pair of jeans, including tax, was $69.62. The following equation can be used to find x, the price of the pair of jeans without the discount. x - 0.1x + 0.07(x - 0.1x) = 69.62 What is the price of the pair of jeans without the discount? $70.58 $72.29 $74.86 $80.02

OpenStudy (anonymous):

0.9x+0.07(0.9x)=69.92 0.9x+0.063x=69.92 0.963x=69.92 x=69.92/0.963=72.6064382139

OpenStudy (anonymous):

so approximately equal to 70.58

OpenStudy (anonymous):

ok thanks cuz that made zero sense